a solution containing 40% alcohol is to be mixed with a solution containing 75% alcohol. How much of each solution should be used to make 100 mL of a mixture that is 50% alcohol?
Answer by checkley77(12844)

.75X+.40(100-X)=.50*100
.75X+40-.40X=50
.35X=50-40
.35X=10
X=10/.35
X=28.57 ML. OF 75% SOLUTION.
100-28.57=71.43 ML. OF 40% SOLUTION.
PROOF:
.75*28.57+40*71.43=.50*100
21.43+28.57=50
50=50